# When Distance Distracts: Representation Distance Bias in BT-Loss for Reward Models

### v1

@@ 3.1 Setup
Base Models. Following standard practice in reward modeling as sequence classifier, we attach a linear score head to a pretrained LLM backbone to output scalar rewards. We experiment with two open-source base models: gemma-2b-it (gemmateam2024gemmaopenmodelsbased) and Llama-3.2-3B-Instruct [^1] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct. These choices evaluate NormBT across distinct backbone families and capacities.
Datasets. We train the reward models on two pairwise preference datasets, yielding a total of four experimental settings (two backbones $\times$ two dataset): (1). Unified-Feedback [^2] https://huggingface.co/datasets/llm-blender/Unified-Feedback. This is a large and diverse collection of pairwise feedback aggregated from multiple open-source datasets. We train on a random subset of 80K preference pairs from Unified-Feedback. (2). Skywork-Reward-Preference-80K-v0.2 (liu2024skyworkrewardbagtricksreward). This is a curated dataset of 80K high-quality preference pairs covering diverse capability and knowledge domains. It is the decontaminated version from Skywork-Reward-Preference-80K-v0.1 (liu2024skyworkrewardbagtricksreward) that removes overlaps with the evaluation benchmark RewardBench, ensuring no contamination between training and evaluation.
Baselines. We compare the performance of NormBT with several loss objectives, including (1) BT Baseline the standard Bradley–Terry loss in Eq 1; (2) BT + Margin (touvron2023llama2openfoundation) which incorporates the ground-truth reward margin $m$ into BT-loss; (3) BT + Margin (outside) (wang2025helpsteer2preferencecomplementingratingspreferences) which varies by introducing $m$ as a scale factor to loss; and (4) BT + label smoothing (liu2024skyworkrewardbagtricksreward) which is a widely used regularization where soft labels replace hard binary targets. The two margin-based variants explicitly modify the prediction-error term by injecting ground-truth reward differences, thereby adjusting update strength according to external signals. Label smoothing also alters the prediction-error term, but by uniformly reducing its magnitude. These methods serve as natural comparisons since they manipulate the error strength, whereas our approach targets representation distance.
For all BT baselines, we conduct an extensive grid search over learning rates and report the best-tuned results. For BT-based variants (margin, margin-out, label smoothing), we search over a range of learning rates centered around the best value found for BT and report the best configuration. Formulation for each baseline and additional training details are included in Appendix B.
@@
(a) Reward models trained from Unified-Feedback (80K)
Reward model Average Chat Chat-Hard Safety Reasoning
@@
BT (baseline) 72.25 95.25 40.35 77.97 75.41
BT + margin 71.23 95.81 37.50 78.65 72.98
BT + margin out 72.53 96.09 38.38 77.57 78.09
BT + label smooth 69.95 93.85 37.72 75.95 72.28
NormBT (ours) 73.57 95.81 39.80 77.97 80.71
@@
BT (baseline) 75.24 95.53 49.89 81.69 71.70
BT + margin 74.15 97.77 47.59 81.28 69.97
BT + margin out 74.84 96.65 45.83 84.05 72.81
BT + label smooth 74.39 95.53 49.78 79.73 72.52
NormBT (ours) 76.96 96.93 49.78 84.19 76.93
(b) Reward models trained from Skywork-Reward-Preference-80K-v0.2
Reward model Average Chat Chat-Hard Safety Reasoning
@@
BT (baseline) 78.63 81.28 73.36 82.43 77.46
NormBT (ours) 80.12 83.80 73.46 82.50 80.71
@@
BT (baseline) 80.31 86.03 78.29 89.86 67.05
NormBT (ours) 81.48 83.80 78.73 88.78 74.60
@@ 3.2 Main Results
Table 1(a) reports the performance of all reward models trained on Unified-Feedback (80K), on both base models. And Table 1(b) presents the results trained on Skywork-Reward-Preference-80K-v0.2. Across all four settings, the NormBT objective consistently outperforms the BT baseline, showing that mitigating representation-distance bias improves reward modeling.
@@
In particular, strong performance gains are observed in the Reasoning category, where NormBT improves accuracy by more than $5\%$ on average. This is consistent with our intuition: preference pairs in the Reasoning category from RewardBench tend to have smaller representation distances (as shown in Figure 2). Under the vanilla BT objective, these pairs receive disproportionately weak updates, even when misranked. By removing this representation-driven effect in parameter updates, NormBT provides stronger updates that are more aligned with prediction errors, leading to the largest performance improvement amongst these pairs.
@@
Analysis. Beyond the results presented in Table 1, we analyze where the performance gains of NormBT originate. Figure 4 shows the subset of RewardBench pairs where the BT Baseline and NormBT predictions disagree. That is, cases where one model is correct while the other is wrong. The x-axis bins response pairs by their representation distance $\|h_{w}-h_{l}\|$, where $h$ are derived from the base model gemma-2b-it, and the y-axis shows the count of these exclusive correct predictions. For ease of interpretation, we divide the range into small, medium, and large distances roughly based on EMA statistics of the embedding distribution; these divisions are intended only for discussion.
The clearest gains of NormBT over BT appear in the small-distance regime. For pairs in this range, BT updates are weak and leave misrankings or ambiguous predictions insufficiently corrected. NormBT resolves this by normalizing the representation distance factor to an average baseline, ensuring prediction error drives the update. Accordingly, the histogram shows a clear concentration of additional correct predictions from NormBT in this region, in line with our intuition. Appendix F shows examples from this region that NormBT predicts correctly while BT doesn’t.
In the medium- and large-distance regimes, results differ slightly across datasets. With Unified-Feedback, NormBT still yields general improvements, while on SKywork-80K both methods perform comparably. This suggests that when the representation distance is large, BT already produces strong updates, and the representation-distance term is less of a bottleneck. By down-weighting pairs from the large-distance range, NormBT maintains parity or modest gains, suggesting that improvements on small-distance pairs do not come at the expense elsewhere.
Overall, these results show that NormBT provides the largest performance gains on pairs with small representation distance, while remaining comparable across other distance ranges.
(a) Unified-Feedback (80K)
(b) Skywork-80K
Figure 4: Comparison on RewardBench pairs where two models disagree. Pairs are binned by representation distance $\|h_{w}-h_{l}\|$ computed from gemma-2B-it. The largest gains for NormBT appear in the small-distance regime, consistent with our analysis that BT under-updates such pairs.
3.3 Ablations and Extensions
To provide a comprehensive evaluation of NormBT, we extend our study to include additional training techniques and scaling variants: (1) LayerNorm, (2) global gradient clipping, (3) active-learning style scaling, and (4) ablations of NormBT and its integration into BT-based variants in Section 3.1. The results are reported in Table 2.
LayerNorm & Gradient Clipping.
A natural question is whether normalizing the representation can mitigate distance-induced imbalance, rather than normalizing the gradient contributions. We therefore apply a LayerNorm before the final reward layer. While this yields mild gains in some categories, it underperforms on fine-grained tasks. The key limitation is that LayerNorm stabilizes embedding scale but does not modify pairwise representation distances, leaving BT gradient dynamics unchanged. Similarly, global gradient clipping avoids extreme gradients, which improves optimization stability but does not address the relative weighting across pairs.
Active-learning Scaling.
Prior active learning works (shen2025revivingclassicsactivereward; feng2025pilafoptimalhumanpreference) prioritize large-distance pairs to improve sample efficiency. Appendix C outlines the key differences between their setting and ours. Nevertheless, for empirical comparison, we adapt the D-opt score formulation (shen2025revivingclassicsactivereward) and extend it with EMA to use as continuous per-pair weights. Although this weighting improves Safety ($77.97\rightarrow 78.99$), it significantly reduces Reasoning performance ($75.41\rightarrow 72.90$), consistent with our view that emphasizing large-distance pairs suppresses fine-grained distinctions. In contrast, NormBT improves Reasoning while maintaining competitive performance in other categories.
NormBT Ablations.
We first ablate on similarity measures used for scaling, replacing the last-token embedding L2-distance with (1) cosine similarity and (2) the L2-distance of average-pooled embeddings. Both alternatives underperform compared to NormBT. This aligns with our theoretical motivation: the last-token embedding is the direct input to the reward head, making its L2 difference an effective proxy for the gradient of the BT score (Eq. 6). Alternative similarity metrics do not reflect this gradient structure. We then remove EMA to show performance degradation as the scale of representation distance drift. Lastly, we apply NormBT on all BT-based variants and observe consistent improvements across all counterparts.
Table 2: Further studies; RMs are trained from gemma-2b-it on Unified-Feedback (80K).
Reward model Average Chat Chat-Hard Safety Reasoning
BT (baseline) 72.25 95.25 40.35 77.97 75.41
BT (LayerNorm) 72.25 94.97 41.34 79.46 73.22
BT (Grad. clip=3.0) 72.25 94.97 41.34 79.46 73.22
BT (Grad. clip=5.0) 72.69 95.53 39.69 77.57 77.96
D-opt Scaling 72.30 96.51 40.79 78.99 72.90
NormBT (No EMA) 67.78 94.97 35.31 75.81 65.04
NormBT (Avg. pool) 69.68 96.09 38.16 73.51 70.97
NormBT (Cos. sim.) 71.88 95.53 39.47 77.16 75.37
NormBT 73.57 95.81 39.80 77.97 80.71
BT + margin 71.23 95.81 37.50 78.65 72.98
NormBT + margin 72.30 95.81 38.82 77.57 77.02
BT + margin out 72.53 96.09 38.38 77.57 78.09
NormBT + margin out 73.36 96.65 38.82 77.16 80.83
BT + label smooth 69.95 93.85 37.72 75.95 72.28
NormBT + label smooth 73.46 95.53 39.25 78.11 80.96
3.4 Downstream RLHF
Following prior works (gao2022scalinglawsrewardmodel; coste2023reward; yang2024regularizing), we conduct BoN sampling on a 1K held-out test set. For each prompt, we sample $n$ candidate responses, score them using the trained RM and select the highest-scoring response. We then use a gold reward model to evaluate the selected responses, and report the gold score averaged over the 1K prompts. This reflects the true quality of the responses selected by RM. We set the number of responses $n$ ranging from 1 to 402 for each prompt. This roughly responds to the KL-divergence of 0 to $5$ from policy model, according to the equation $\text{KL}_{\text{{BoN}}}=\ln n-\frac{n-1}{n}$. The gold score model333reward-model-Mistral-7B-instruct-Unified-Feedback is a 7B RM finetuned on the entire Unified-Feedback dataset, following yang2024regularizing.
The results are reported in Figure 5. NormBT consistently achieves higher gold scores, outperforming all BT baselines. Moreover, applying NormBT on top of each BT variant yields further improvements, most notably for label smoothing. This demonstrates that NormBT complements and strengthens existing approaches in ways that carry over to downstream RLHF applications.
@@ 3.4 Downstream RLHF
Figure 5: Best-of-N Selection. Higher gold score indicates higher quality of the selected responses. NormBT consistently outperforms all BT baselines, as well as BT counterparts in ablation studies.
@@ 4 Related Works
Reward modeling is central to reinforcement learning from human feedback (RLHF), aligning LLMs with human preferences (christiano2017deep; stiennon2022learningsummarizehumanfeedback; bai2022traininghelpfulharmlessassistant; dong2024rlhfworkflowrewardmodeling). Reward models can be broadly classified into paradigms. The first is discriminative RM, typically with a linear score head trained via a Bradley-Terry (BT) style objective (wang2025helpsteer2preferencecomplementingratingspreferences; liu2024skyworkrewardbagtricksreward; yang2024regularizing). Another line of research focuses on generative RMs, aiming to leverage the model’s generation ability to produce rationales, critiques, or verifier signals (mahan2024generativerewardmodels; zhang2025generativeverifiersrewardmodeling; chen2025rmr1rewardmodelingreasoning; zhu2025retrievalaugmentedprocessrewardmodel; chen2025reasongrmenhancinggenerativereward). Beyond this dichotomy, several branches enrich the supervision signal, such as providing fine-grained feedback (wu2023finegrainedhumanfeedbackgives), incorporating multi-objectives (wang2024interpretablepreferencesmultiobjectivereward), and providing awareness of distribution or uncertainty (dorka2024quantileregressiondistributionalreward; sun2025probabilisticuncertainrewardmodel). In this work, we focus on reward models in the discriminative, BT-based paradigm. While Direct alignment methods (rafailov2024directpreferenceoptimizationlanguage; meng2024simposimplepreferenceoptimization; gupta2025alphaporewardshapematters) bypass explicit reward modeling, they implicitly rely on similar BT-style pairwise signals. Our work provides a principled analysis of how pairwise signals translate into parameter updates.
@@
Despite the success of reward models in RLHF, a broad literature targets complementary weaknesses in training and deployment. For instance,  liu2024reward incorporates ties into BT models, while  coste2023reward identifies the issue of overoptimization.  yang2024regularizing shows that regularizing hidden states helps generalization against distribution shifts. Data-centric improvements curate higher-quality comparisons to enhance reward modeling (liu2024rm; cui2023ultrafeedback; liu2025skyworkrewardv2scalingpreferencedata). A parallel line of work improves evaluation protocols, aiming to better capture the practical utility of reward models (liu2024rmbenchbenchmarkingrewardmodels; frick2024evaluaterewardmodelsrlhf; malik2025rewardbench2advancingreward). Our study is orthogonal to these directions, where we address a structural limitation in BT-style updates without modifying data or model architecture.  hong2025robustnessrewardmodelslanguage also analyzes representation distance in RMs and introduces zero-centered reward regularization. Our work analyzes how each pair directly contributes to model updates and proposes to modify the BT gradients explicitly. Finally, theoretical and diagnostic studies (razin2025makesrewardmodelgood; sun2025rethinkingbradleyterrymodelspreferencebased) analyze fundamental drawbacks of reward models. Our study is complementary to these works in exploring the limitations as well as further possibilities of the BT objective.

### v2

@@ 3 Experiments
We conduct comprehensive experiments to evaluate NormBT against BT baselines, compare downstream performance in Section 3.3, followed by ablation studies in Section 3.4 and analysis of performance gains in Section 3.5.
@@ 3.1 Setup
Base Models. Following standard practice in reward modeling as sequence classifier, we attach a linear score head to a pretrained LLM backbone to output scalar rewards. We experiment with two open-source base models: gemma-2b-it (Team et al., 2024) and Llama-3.2-3B-Instruct111meta-llama/Llama-3.2-3B-Instruct. These choices evaluate NormBT across distinct backbone families and capacities.
Datasets. We train reward models on two pairwise preference datasets, yielding a total of four experimental settings (two backbones $\times$ two dataset): (1). Unified-Feedback222llm-blender/Unified-Feedback. This is a large and diverse collection of pairwise feedback aggregated from multiple open-source datasets. We use a random subset of 80K preference pairs from Unified-Feedback. (2). Skywork-Reward-Preference-80K-v0.2 (Liu et al., 2024a). This is a curated dataset of 80K high-quality preference pairs covering diverse capability and knowledge domains. It is the decontaminated version from Skywork-Reward-Preference-80K-v0.1 (Liu et al., 2024a) to remove overlaps with the evaluation benchmark RewardBench, ensuring no contamination between training and evaluation.
Baselines. We compare NormBT against several reward model loss objectives:
  1. 1.
    BT Baseline: standard Bradley–Terry loss in Eq. 1.
  2. 2.
    BT + Margin (Touvron et al., 2023): introduces the ground-truth reward margin $m$ into the BT loss.
  3. 3.
    BT + Margin (outside) (Wang et al., 2025): variant that introduces margin $m$ as a scale factor to the loss.
  4. 4.
    BT + label smoothing (Liu et al., 2024a): replaces hard binary targets with soft labels as a regularizer.
The two margin-based variants explicitly modify the prediction-error term by injecting ground-truth reward differences, thereby adjusting update strength according to external signals. Label smoothing also alters the prediction-error term, but by uniformly reducing its magnitude. These methods serve as natural comparisons since they manipulate the error strength, whereas our approach targets representation distance. Detailed formulation and comparison with NormBT are included in Appendix B.
For all BT baselines, we conduct an extensive grid search over learning rates and report the best-tuned results. Additional training details are included in Appendix B.
@@
(a) Models trained from Unified-Feedback (80K)
Reward model Chat Chat-Hard Safety Reasoning Average
@@
BT (baseline) 95.25 40.35 77.97 75.41 72.25
BT + margin 95.81 37.50 78.65 72.98 71.23
BT + margin out 96.09 38.38 77.57 78.09 72.53
BT + label smooth 93.85 37.72 75.95 72.28 69.95
NormBT (ours) 95.81 39.80 77.97 80.71 73.57
@@
BT (baseline) 95.53 49.89 81.69 71.70 75.24
BT + margin 97.77 47.59 81.28 69.97 74.15
BT + margin out 96.65 45.83 84.05 72.81 74.84
BT + label smooth 95.53 49.78 79.73 72.52 74.39
NormBT (ours) 96.93 49.78 84.19 76.93 76.96
(b) Models trained from Skywork-Reward-Preference-80K-v0.2
Reward model Chat Chat-Hard Safety Reasoning Average
@@
BT (baseline) 81.28 73.36 82.43 77.46 78.63
NormBT (ours) 83.80 73.46 82.50 80.71 80.12
@@
BT (baseline) 86.03 78.29 89.86 67.05 80.31
NormBT (ours) 83.80 78.73 88.78 74.60 81.48
@@ 3.2 Main Results
Table 1(a) reports the performance of all reward models trained on Unified-Feedback (80K) and the two base models. And Table 1(b) presents the results trained on Skywork-Reward-Preference-80K-v0.2. Across all four settings, the NormBT objective consistently outperforms the BT baseline, showing that mitigating representation-distance bias improves reward modeling.
@@
In particular, strong performance gains are observed in the Reasoning category, where NormBT improves accuracy by more than $5\%$ on average. This is consistent with our analysis: Reasoning pairs tend to have smaller representation distances compared to other categories (Figure 2(b)). Under BT, the small-distance region receives inherently weak updates, even when misranked. NormBT explicitly alleviates this by upweighting learning signals for small-distance regions and reducing over-emphasis on large-distance regions. By removing this representation-driven effect in parameter updates, NormBT provides stronger updates that are more aligned with prediction errors. As a result, the largest gains naturally appear in categories dominated by fine-grained distinctions, which are otherwise suppressed.
@@
3.3 Downstream Performance
@@
Figure 3: Best-of-N Selection. Higher gold score indicates higher quality of the selected responses. (a) NormBT consistently outperforms all BT baselines, as well as (b) all BT-based counterparts when applied on top of each BT variant.
We evaluate the reward models from Section 3.1 trained on gemma-2b-it and Unified-Feedback (80K). We further applied NormBT on top of each BT baselines to examine the effectiveness of our method on downstream performance.
Following prior works (Gao et al., 2022; Coste et al., 2023; Yang et al., 2024), we conduct BoN sampling on a 1K held-out test set. For each prompt, we sample $n$ candidate responses, score them using the trained RM and select the highest-scoring response. We then use a gold reward model to evaluate the selected responses, and report the gold score averaged over the 1K prompts. This reflects the true quality of the responses selected by RM. We set the number of responses $n$ ranging from 1 to 402 for each prompt. This roughly responds to the KL-divergence of 0 to $5$ from policy model, according to the equation $\text{KL}_{\text{{BoN}}}=\ln n-\frac{n-1}{n}$. The gold score model333reward-model-Mistral-7B-instruct-Unified-Feedback is a 7B RM finetuned on the entire Unified-Feedback dataset, following (Yang et al., 2024).
The results are reported in Figure 3. NormBT consistently achieves higher gold scores, outperforming all BT baselines. Moreover, applying NormBT on top of each BT variant yields further improvements , showing that balancing the effect of representation distance on learning signals is beneficial to standalone BT objective. The performance gains is most pronouced for label smoothing: smoothing uniformly damps gradients and can further weaken already-small updates for small-distance pairs, whereas NormBT compensates by upweighting such pairs, leading to a clear boost in BoN gold score. These results demonstrate that NormBT complements and strengthens existing approaches in ways that carry over to downstream RLHF applications.
3.4 Ablation Studies
We compare NormBT against standard training techniques and scaling variants: (1) LayerNorm, (2) global gradient clipping, (3) ablations of NormBT, and (4) its integration into BT-based variants. The results are reported in Table 2.
LayerNorm & Gradient Clipping. A natural question is whether uniformly normalizing the representation can mitigate distance-induced imbalance, rather than reweighting the gradient contributions. We therefore apply a LayerNorm before the final reward layer. While this yields mild gains in some categories, it underperforms on fine-grained tasks. The key limitation is that LayerNorm stabilizes embedding scale but does not modify the disparity in representation distances, leaving BT gradient dynamics unchanged. Similarly, global gradient clipping avoids extreme gradients, which improves optimization stability but does not address the relative weighting across pairs.
Ablations. We first ablate on similarity measures used for scaling, replacing the last-token embedding L2-distance with (1) cosine similarity and (2) the L2-distance of average-pooled embeddings. Both alternatives underperform compared to NormBT. This aligns with our theoretical motivation: the last-token embedding is the direct input to the reward head, making its L2 difference an effective proxy for the gradient of the BT score (Eq. 6). Alternative similarity metrics do not reflect this gradient structure. We also remove EMA and observe performance degradation, as the scale of representation distance drift during training (Appendix G). Lastly, we apply NormBT on all BT-based variants and observe consistent improvements across all counterparts.
Table 2: Ablation Studies. All reward models are trained from gemma-2b-it on Unified-Feedback (80K).
Reward model Chat Chat-Hard Safety Reasoning Average
BT (baseline) 95.25 40.35 77.97 75.41 72.25
BT (LayerNorm) 94.97 41.34 79.46 73.22 72.25
BT (Grad. clip=3.0) 94.97 41.34 79.46 73.22 72.25
BT (Grad. clip=5.0) 95.53 39.69 77.57 77.96 72.69
NormBT (No EMA) 94.97 35.31 75.81 65.04 67.78
NormBT (Avg. pool) 96.09 38.16 73.51 70.97 69.68
NormBT (Cos. sim.) 95.53 39.47 77.16 75.37 71.88
NormBT 95.81 39.80 77.97 80.71 73.57
BT + margin 95.81 37.50 78.65 72.98 71.23
NormBT + margin 95.81 38.82 77.57 77.02 72.30
BT + margin out 96.09 38.38 77.57 78.09 72.53
NormBT + margin out 96.65 38.82 77.16 80.83 73.36
BT + label smooth 93.85 37.72 75.95 72.28 69.95
NormBT + label smooth 95.53 39.25 78.11 80.96 73.46
3.5 Analysis
Beyond the evaluation results presented, we analyze where the performance gains of NormBT originate. Figure 4 shows the subset of RewardBench pairs where the BT Baseline and NormBT predictions disagree. That is, cases where one model is correct while the other is wrong. The x-axis bins response pairs by their representation distance $\|h_{w}-h_{l}\|$, where $h$ are derived from the base model gemma-2b-it, and the y-axis shows the count of these exclusive correct predictions. For ease of interpretation, we divide the range into small, medium, and large distances roughly based on EMA statistics of the embedding distribution; these divisions are intended only for discussion.
The clearest gains of NormBT over BT appear in the small-distance regime. For pairs in this range, BT updates are weak and leave misrankings or ambiguous predictions insufficiently corrected. NormBT resolves this by normalizing the representation distance factor to an average baseline, ensuring prediction error drives the update. Accordingly, the histogram shows a clear concentration of additional correct predictions from NormBT in this region, in line with our intuition. Appendix D provides examples from this region that NormBT predicts correctly while BT doesn’t.
In the medium- and large-distance regimes, NormBT still yields general improvements. This is consistent with our analysis: when representation distance is large, BT already produces strong updates, so the distance term is less of a bottleneck. We do observe small degradations relative to the best baseline, for example, $78.65\rightarrow 77.97$ on Safety for gemma-2b-it trained on Unified-Feedback (Table 1), suggesting that down-weighting can slightly affect large-distance regions. Overall, NormBT maintains parity or modest gains, suggesting that improvements on small-distance pairs come at little expense elsewhere.
Figure 4: Comparison on RewardBench pairs where two models disagree (Unified-Feedback, 80K). Pairs are binned by representation distance $\|h_{w}-h_{l}\|$ computed from gemma-2B-it. The largest gains for NormBT appear in the small-distance regime, consistent with our analysis that BT under-updates such pairs.
@@ 4 Related Works
Reward modeling is central to reinforcement learning from human feedback (RLHF), aligning LLMs with human preferences (Christiano et al., 2017; Stiennon et al., 2022; Bai et al., 2022; Dong et al., 2024). Reward models can be broadly classified into paradigms. The first is discriminative RM, typically with a linear score head trained via a Bradley-Terry (BT) style objective (Wang et al., 2025; Liu et al., 2024a; Yang et al., 2024). Another line of research focuses on generative RMs, aiming to leverage the model’s generation ability to produce rationales, critiques, or verifier signals (Mahan et al., 2024; Zhang et al., 2025; Chen et al., 2025b; Zhu et al., 2025; Chen et al., 2025a). Beyond this dichotomy, several branches enrich the supervision signal, such as providing fine-grained feedback (Wu et al., 2023), incorporating multi-objectives (Wang et al., 2024), and providing awareness of distribution or uncertainty (Dorka, 2024; Sun et al., 2025b). In this work, we focus on reward models in the discriminative, BT-based paradigm. While Direct alignment methods (Rafailov et al., 2024; Meng et al., 2024; Gupta et al., 2025) bypass explicit reward modeling, they implicitly rely on similar BT-style pairwise signals. Our work provides a principled analysis of how pairwise signals translate into parameter updates.
@@
Despite the success of reward models in RLHF, a broad literature targets complementary weaknesses in training and deployment. For instance,  (Liu et al., 2024b) incorporates ties into BT models, while  (Coste et al., 2023) identifies the issue of overoptimization.  (Yang et al., 2024) shows that regularizing hidden states helps generalization against distribution shifts. Data-centric improvements curate higher-quality comparisons to enhance reward modeling (Liu et al., 2024c; Cui et al., 2023; Liu et al., 2025; Shen et al., 2025). A parallel line of work improves evaluation protocols, aiming to better capture the practical utility of reward models (Liu et al., 2024d; Frick et al., 2024; Malik et al., 2025). Our study is orthogonal to these directions, where we address a structural limitation in BT-style updates without modifying data or model architecture.  (Hong et al., 2025) also analyzes representation distance in RMs and introduces zero-centered reward regularization. Our work analyzes how each pair directly contributes to model updates and proposes to modify the BT gradients explicitly. Finally, theoretical and diagnostic studies (Razin et al., 2025; Sun et al., 2025a) analyze fundamental drawbacks of reward models. Our study is complementary to these works in exploring the limitations as well as further possibilities of the BT objective.